OpenStudy (anonymous):

Olivia has $250 to buy presents for her 3 sisters and 2 nephews. She wants to spend $10 more on each sister than she spends on each nephew. Let s = the amount she spends on each sister, and n = the amount she spends on each nephew. How much (in dollars) does she spend on each sister?

OpenStudy (anonymous):

So we can write some equations to give us a system that can be solved. You an always write an equation for the total: 250 = 3s + 2n And we can relate the price of a sister's gift to that of a nephew: s = n + 10 Then we take these two equations and substitute: 250 = 3(n + 10) + 2n 250 = 3n + 30 + 2n 220 = 5n n = 44 To find the sister's price, we simply substitute into the price relationship: s = n + 10 s = 44 + 10 s = 54 So $54 will be spent on each sister

OpenStudy (anonymous):

THANK YOU!